# Continual Learning: Forget-free Winning Subnetworks for Video Representations

### 19 Dec 2023 v1

@@ Abstract
Inspired by the Regularized Lottery Ticket Hypothesis (RLTH), which highlights the presence of competitive subnetworks within dense networks for continual learning tasks, we introduce Winning Subnetworks (WSN). This approach utilizes reused weights in dense networks to enhance learning in Task Incremental Learning (TIL) scenarios. To mitigate overfitting in Few-Shot Class Incremental Learning (FSCIL), we have developed WSN variants referred to as the Soft subnetwork (SoftNet). Furthermore, addressing WSN’s limitation of sparse reused weights in Video Incremental Learning (VIL), we propose the Fourier Subneural Operator (FSO). The FSO, operating in Fourier space, adaptively and compactly encodes videos, discovering reusable subnetworks with diverse bandwidths. We have applied FSO’s Fourier representations to various continual learning contexts, including VIL, TIL, and FSCIL. Our extensive experiments across these scenarios demonstrate FSO’s remarkable efficacy in continual learning, significantly enhancing task performance at various convolutional representational levels: it boosts performance in the higher layers for TIL and FSCIL and the lower layers for VIL.
@@ 1 Introduction
Continual Learning (CL), also known as Lifelong Learning [82, 66, 96, 24], is a learning paradigm where a series of tasks are learned sequentially. The principle objective of continual learning is to replicate human cognition, characterized by the ability to learn new concepts or skills incrementally throughout one’s lifespan. An optimal continual learning system could facilitate a positive forward and backward transfer, leveraging the knowledge gained from previous tasks to solve new ones while also updating its understanding of previous tasks with the new knowledge. However, building successful continual learning systems is challenging due to the occurrence of catastrophic forgetting or catastrophic interference [55], a phenomenon where the performance of the model on previous tasks significantly deteriorates when it learns new tasks. This can make it challenging to retain the knowledge acquired from previous tasks, ultimately leading to a decline in overall performance. To tackle the catastrophic forgetting problem in continual learning, numerous approaches have been proposed, which can be broadly classified as follows: (1) Regularization-based methods [37, 4, 31, 83, 59] aim to keep the learned information of past tasks during continual training aided by sophisticatedly designed regularization terms, (2) Rehearsal-based methods [64, 65, 5, 6, 67, 68] utilize a set of real or synthesized data from the previous tasks and revisit them, and (3) Architecture-based methods [52, 70, 44, 89, 33, 34] propose to minimize the inter-task interference via newly designed architectural components.
@@
Despite the remarkable success of recent works on rehearsal- and architecture-based continual learning, most current methods request external memory as new tasks arrive, making the model difficult to scale to larger and more complex tasks. Rehearsal-based CL requires additional storage to store the replay buffer or generative models, and architecture-based methods leverage additional model capacity to account for new tasks. These trends lead to an essential question: how can we build a memory-efficient CL model that does not exceed the backbone network’s capacity or even requires a much smaller capacity? Several studies have shown that deep neural networks are over-parameterized [18, 23, 43] and thus removing redundant/unnecessary weights can achieve on-par or even better performance than the original dense network. More recently, Lottery Ticket Hypothesis (LTH) [20] demonstrates the existence of sparse subnetworks, named winning tickets, that preserve the performance of a dense network. However, searching for optimal winning tickets during continual learning with iterative pruning methods requires repetitive pruning and retraining for each arriving task, which could be more practical.
@@
To tackle the issues of external replay buffer, capacity, and over-fitting, we suggest a novel CL method, which finds the high-performing Winning SubNetwork referred to as WSN [33] given tasks without the need for retraining and rewinding, as shown in Figure 1 (d). Also, we set previous pruning-based CL approaches [52, 89] (see Figure 1 (a)) to baselines of architectures, which obtain task-specific subnetworks given a pre-trained backbone network. Our WSN incrementally learns model weights and task-adaptive binary masks (the subnetworks) within the neural network. To allow the forward transfer when a model learns on a new task, we reuse the learned subnetwork weights for the previous tasks, however selectively, as opposed to using all the weights [53] (see Figure 1 (b)), that may lead to biased transfer. Further, the WSN eliminates the threat of catastrophic forgetting during continual learning by freezing the subnetwork weights for the previous tasks and does not suffer from the negative transfer, unlike [95] (see Figure 1 (c)), whose subnetwork weights for the previous tasks can be updated when training on the new sessions. Moreover, we observed that subnetworks could overfit to limited task data, potentially reducing their effectiveness on new tasks or datasets, such as in Few-Shot Class Incremental Learning (FSCIL). To address this, we adopted the Regularized Lottery Ticket Hypothesis (RLTH) [34]. This led to the discovery of regularized subnetworks characterized by smoother (soft) masks referred to as Soft-SubNetwork (SoftNet), enhancing their adaptability and performance.
@@
While these conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting, they are unsuited for sequential complex Video Continual Learning (VCL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complicated oscillatory signals) in frequency space. To achieve this, we define Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as Task Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO.
@@
  • We have expanded the FSO of Fourier representations to encompass a variety of continual learning scenarios: Video Incremental Learning (VIL), Task Incremental Learning (TIL), and Few-Shot Class Incremental Learning (FSCIL). In our evaluations, the proposed FSO demonstrates superior performance compared to architecture-based continual learning models, such as WSN and SoftNet in TIL, VIL, and FSCIL, respectively, thereby underscoring its exceptional representational power.
@@ 2 Related Works
Continual Learning [55, 82, 41, 45], also known as lifelong learning, is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is regularization-based methods [37, 4, 31, 83, 59], which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. Rehearsal-based approaches [64, 5, 6, 67, 17, 79, 68] store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some methods in this line of work [73, 2] accommodate the generative model to construct the pseudo-rehearsals for previous tasks. Architecture-based approaches [52, 70, 44, 89, 33, 34] use the additional capacity to expand [91, 95], dynamic representation [92, 74] or isolate [66] model parameters, preserving learned knowledge and preventing forgetting. Both rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters [89] or the replay buffers. Prompt-based learning, an emerging transfer learning technique in the field of natural language processing (NLP), harnesses a fixed function to condition the model. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Notably, while L2P [88] stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt [87] introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox [19], S-Prompt [86], CODA-P [77], ConStruct-VL [76], ST-Prompt [61], and LGCL [35]. Nonetheless, the central focus of this study is to pinpoint the most optimal winning tickets for convolutional operators in a wide array of continual learning scenarios, both in real and Fourier space.
@@
Architecture-based Continual Learning. While most works aim to increase the performance of continual learners by adding memory, some researchers have focused on building memory and computationally efficient continual learners by using pruning-based constraints. CLNP [21] is one example of a method that selects important neurons for a given task using $\ell_{1}$ regularization to induce sparsity and freezes them to maintain performance. Neurons that are not selected are reinitialized for future task training. Another method, Piggyback [52], trains task-specific binary masks on the weights given a pre-trained model. However, this method does not allow for knowledge transfer among tasks, and its performance highly depends on the quality of the backbone model. HAT [70] proposes task-specific learnable attention vectors to identify important weights per task. The masks are formulated to layerwise cumulative attention vectors during continual learning. A recent method, LL-Tickets [12], shows a sparse subnetwork called lifelong tickets that performs well on all tasks during continual learning. The method searches for more prominent tickets from current ones if the obtained tickets cannot sufficiently learn the new task while maintaining performance on past tasks. However, LL-Tickets require external data to maximize knowledge distillation with learned models for prior tasks, and the ticket expansion process involves retraining and pruning steps. WSN [33] is another method that jointly learns the model and task-adaptive binary masks on subnetworks associated with each task. It attempts to select a small set of weights activated (winning ticket) by reusing weights of the prior subnetworks in Task Incremental Learning (TIL). We also investigate the performance of Soft-Subnetworks (SoftNet) [34]. We select sub-networks (winning tickets) to obtain base session knowledge and optimize the regularized sub-network parameters for new sessions in Few-Shot Class Incremental Learning (FSCIL) settings.
@@
Neural Implicit Representation (NIR) [56] are neural network architectures for parameterizing continuous, differentiable signals. Based on coordinate information, they provide a way to represent complex, high-dimensional data with a small set of learnable parameters that can be used for various tasks such as image reconstruction [75, 80], shape regression [13, 60], and 3D view synthesis  [58, 69]. Instead of using coordinate-based methods, NeRV [10] proposes an image-wise implicit representation that takes frame indices as inputs, enabling fast and accurate video compression. NeRV has inspired further improvements in video regression by CNeRV [8], DNeRV [25], E-NeRV [46], and NIRVANA [51], and HNeRV [9]. A few recent works have explored video continual learning (VCL) scenarios for the NIR. To tackle non-physical environments, Continual Predictive Learning (CPL) [7] learns a mixture world model via predictive experience replay and performs test-time adaptation using non-parametric task inference. PIVOT [85] leverages the past knowledge present in pre-trained models from the image domain to reduce the number of trainable parameters and mitigate forgetting. CPL needs memory to replay, while PIVOT needs pre-training and fine-tuning steps. In contrast, along with the conventional progressive training techniques [66, 16], we use WSN as baselines, which utilizes the Lottery Ticket Hypothesis (LTH) to identify an adaptive substructure within the dense networks that are tailored to the specific video input index. However, WSN is inappropriate for sequential complex video compilation since it reuses a few adaptive but sparse learnable parameters. To overcome the weakness of WSN, We proposed a novel Fourier Subneural Operator (FSO) for representing complex video in Fourier space [47, 48, 38, 84]. We have expanded the FSO of Fourier representations to encompass a variety of continual learning architectures and scenarios to validate its effectivness.
@@ 3 Winning SubNetworks in Fourier Space
In this section, we present our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[33]. The WSN searches for the task-adaptive winning tickets and updates only the weights that have not been trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [34], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting, however, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, VIL, and FSCIL, to demonstrate the effectiveness of FSO.
@@ 5 Results of Task-Incremental Learning
Table I: Performance comparisons of the proposed method and other state-of-the-art including baselines - PackNet [53] and SupSup [89] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [17]. The best results are highlighted in bold. Also, $\dagger$    denotes results reported from [17].

### 04 Jan 2024 v2

@@ Abstract
Inspired by the Lottery Ticket Hypothesis (LTH), which highlights the existence of efficient subnetworks within larger, dense networks, a high-performing Winning Subnetwork (WSN) in terms of task performance under appropriate sparsity conditions is considered for various continual learning tasks. It leverages pre-existing weights from dense networks to achieve efficient learning in Task Incremental Learning (TIL) scenarios. In Few-Shot Class Incremental Learning (FSCIL), a variation of WSN referred to as the Soft subnetwork (SoftNet) is designed to prevent overfitting when the data samples are scarce. Furthermore, the sparse reuse of WSN weights is considered for Video Incremental Learning (VIL). The use of Fourier Subneural Operator (FSO) within WSN is considered. It enables compact encoding of videos and identifies reusable subnetworks across varying bandwidths. We have integrated FSO into different architectural frameworks for continual learning, including VIL, TIL, and FSCIL. Our comprehensive experiments demonstrate FSO’s effectiveness, significantly improving task performance at various convolutional representational levels. Specifically, FSO enhances higher-layer performance in TIL and FSCIL and lower-layer performance in VIL.
@@ 1 Introduction
Continual Learning (CL), also known as Lifelong Learning [83, 67, 97, 24], is a learning paradigm where a series of tasks are learned sequentially. The principle objective of continual learning is to replicate human cognition, characterized by the ability to learn new concepts or skills incrementally throughout one’s lifespan. An optimal continual learning algorithm could facilitate a positive forward and backward transfer, leveraging the knowledge gained from previous tasks to solve new ones while also updating its understanding of previous tasks with the new knowledge. However, building successful continual learning algorithms is challenging due to the occurrence of catastrophic forgetting or catastrophic interference [56], a phenomenon where the performance of the model on previous tasks significantly deteriorates when learning new tasks. This can make it challenging to retain the knowledge acquired from previous tasks, ultimately leading to a decline in overall performance. To tackle the catastrophic forgetting problem in continual learning, numerous approaches have been proposed, which can be broadly classified as follows: (1) Regularization-based methods [38, 4, 31, 84, 60] aim to keep the learned information of past tasks during continual training aided by sophisticatedly designed regularization terms, (2) Rehearsal-based methods [65, 66, 5, 6, 68, 69] utilize a set of real or synthesized data from the previous tasks and revisit them, and (3) Architecture-based methods [53, 71, 45, 90, 34, 35] propose to minimize the inter-task interference via newly designed architectural components.
@@
Despite the remarkable success of recent works on rehearsal- and architecture-based continual learning, most current methods request external memory as new tasks arrive, making the model difficult to scale to larger and more complex tasks. Rehearsal-based CL requires additional storage to store the replay buffer or generative models, and architecture-based methods leverage additional model capacity to account for new tasks. These trends lead to an essential question: how can we build a memory-efficient CL model that does not exceed the backbone network’s capacity or even requires a much smaller capacity? Several studies have shown that deep neural networks are over-parameterized [18, 23, 44] and thus removing redundant/unnecessary weights can achieve on-par or even better performance than the original dense network. More recently, Lottery Ticket Hypothesis (LTH) [20] demonstrates the existence of sparse subnetworks, named winning tickets, that preserve the performance of a dense network. However, searching for optimal winning tickets during continual learning with iterative pruning methods requires repetitive pruning and retraining for each arriving task, which could be more impractical.
@@
To tackle the issues of external replay buffer and capacity, we suggest a novel CL method, which finds the high-performing Winning SubNetwork referred to as WSN [34] given tasks without the need for retraining and rewinding, as shown in Figure 1 (d). Also, we set previous pruning-based CL approaches [53, 90] (see Figure 1 (a)) to baselines of architectures, which obtain task-specific subnetworks given a pre-trained backbone network. Our WSN incrementally learns model weights and task-adaptive binary masks (the subnetworks) within the neural network. To allow the forward transfer when a model learns on a new task, we reuse the learned subnetwork weights for the previous tasks, however selectively, as opposed to using all the weights [54] (see Figure 1 (b)), that may lead to biased transfer. Further, the WSN eliminates the threat of catastrophic forgetting during continual learning by freezing the subnetwork weights for the previous tasks and does not suffer from the negative transfer, unlike [96] (see Figure 1 (c)), that subnetwork weights for the previous tasks can be updated when training on the new sessions. Moreover, we observed that subnetworks could overfit limited task sample data, potentially reducing their effectiveness on new tasks or datasets, such as in Few-Shot Class Incremental Learning (FSCIL). To address the overfitting issue, we adopted the Regularized Lottery Ticket Hypothesis (RLTH) [35]. This led to the discovery of regularized subnetworks characterized by smoother (soft) masks referred to as Soft-SubNetwork (SoftNet), enhancing their adaptability and performance.
@@
These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting or to alleviate overfitting. However, they are unsuited for sequential complex Video Incremental Learning (VIL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complex signals) in frequency space. To achieve this, we define Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. Given a backbone and FSO architecture, our method continuously learns to identify input-adaptive sub-modules in Fourier space and encode videos in each sequential training session. We apply FSO to various architectures accompanied by continual learning scenarios, such as Task Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO representations.
@@
  • We have applied the FSO to various architectures used by a variety of continual learning scenarios: Video Incremental Learning (VIL), Task Incremental Learning (TIL), and Few-Shot Class Incremental Learning (FSCIL). In our evaluations, the proposed FSO performs better than architecture-based continual learning models, such as WSN and SoftNet, in TIL, VIL, and FSCIL, respectively, underscoring its exceptional representational power.
@@ 2 Related Works
Continual Learning [56, 83, 42, 46], also known as lifelong learning, is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is regularization-based methods [38, 4, 31, 84, 60], which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. Rehearsal-based approaches [65, 5, 6, 68, 17, 80, 69] store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some methods in this line of work [74, 2] accommodate the generative model to construct the pseudo-rehearsals for previous tasks. Architecture-based approaches [53, 71, 45, 90, 34, 35] use the additional capacity to expand [92, 96], dynamic representation [93, 75] or isolate [67] model parameters, preserving learned knowledge and preventing forgetting. Both rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters [90] or the replay buffers. Recently, Prompt-based learning, an emerging transfer learning technique in natural language processing (NLP), harnesses a fixed function of pre-trained Transformer models. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Notably, while L2P [89] stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt [88] introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox [19], S-Prompt [87], CODA-P [78], ConStruct-VL [77], ST-Prompt [62], and LGCL [36]. All previous learning methods depend mainly on continual representations in real space. However, the central focus of this study is to pinpoint the most optimal winning ticket representations for convolutional operators in three continual learning scenarios in Fourier space.
@@
Architecture-based Continual Learning. While most works aim to increase the performance of continual learners by adding memory, some researchers have focused on building memory and computationally efficient continual learners by using pruning-based constraints. CLNP [21] is one example of a method that selects important neurons for a given task using $\ell_{1}$ regularization to induce sparsity and freezes them to maintain performance. Neurons that are not selected are reinitialized for future task training. Another method, Piggyback [53], trains task-specific binary masks on the weights given a pre-trained model. However, this method does not allow for knowledge transfer among tasks, and its performance highly depends on the quality of the backbone model. HAT [71] proposes task-specific learnable attention vectors to identify significant weights per task. The masks are formulated to layerwise cumulative attention vectors during continual learning. A recent method, LL-Tickets [12], shows a sparse subnetwork called lifelong tickets that performs well on all tasks during continual learning. The method searches for more prominent tickets from current ones if the obtained tickets cannot sufficiently learn the new task while maintaining performance on past tasks. However, LL-Tickets require external data to maximize knowledge distillation with learned models for prior tasks, and the ticket expansion process involves retraining and pruning steps. WSN [34] is a new method that jointly learns the model and task-adaptive binary masks on subnetworks associated with each task, which can be extended to Fourier spaces. It attempts to select a small set of weights activated (winning ticket) by reusing weights of the prior subnetworks in Task Incremental Learning (TIL). Also, in Few-Shot Class Incremental Learning (FSCIL), we investigate the performance of Soft-Subnetworks (SoftNet) [35] as another variant of WSN: we select sub-networks (winning tickets) to obtain base session knowledge and optimize the regularized sub-network parameters to alleviate overfitting few samples for new sessions.
@@
Neural Implicit Representation (NIR) [57] are neural network architectures for parameterizing continuous, differentiable signals. Based on coordinate information, they provide a way to represent complex, high-dimensional data with a small set of learnable parameters that can be used for various tasks such as image reconstruction [76, 81], shape regression [13, 61], and 3D view synthesis  [59, 70]. Instead of using coordinate-based methods, NeRV [10] proposes an image-wise implicit representation that takes frame indices as inputs, enabling fast and accurate video compression. NeRV has inspired further improvements in video regression by CNeRV [8], DNeRV [25], E-NeRV [47], and NIRVANA [52], and HNeRV [9]. A few recent works have explored video continual learning (VCL) scenarios for the NIR. To tackle non-physical environments, Continual Predictive Learning (CPL) [7] learns a mixture world model via predictive experience replay and performs test-time adaptation using non-parametric task inference. PIVOT [86] leverages the past knowledge present in pre-trained models from the image domain to reduce the number of trainable parameters and mitigate forgetting. CPL needs memory to replay, while PIVOT needs pre-training and fine-tuning steps. In contrast, along with the conventional progressive training techniques [67, 16], we use WSN as baselines, which utilizes the Lottery Ticket Hypothesis (LTH) to identify an adaptive substructure within the dense networks that are tailored to the specific video input index. However, WSN is inappropriate for sequential complex video compilation since it reuses a few adaptive but sparse learnable parameters. To overcome the weakness of WSN, We proposed a novel Fourier Subneural Operator (FSO) [33] for representing complex video in Fourier space [48, 49, 39, 85]. We have expanded the FSO of Fourier representations to encompass a variety of continual learning architectures and scenarios to validate its effectivness.
@@ 3 Winning SubNetworks in Fourier Space
This section presents our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[34] and introduces the Fourier Subnetwork Operator (FSO, see Figure 3 and Figure 4) for better video representations. Then, we depict how we apply FSO to various architectures used in three continual learning scenarios.
3.1 WSN & Fourier Subnetwork Operator (FSO)
The WSN searches for the task-adaptive winning tickets and updates only the weights not trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [35], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting. However, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subnetwork Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, VIL, and FSCIL, to demonstrate the effectiveness of FSO.
@@ 5 Results of Task-Incremental Learning
Table I: Performance comparisons of the proposed method and other baselines - PackNet [54] and SupSup [90] - on various benchmark datasets. We report the mean and standard deviation of the average accuracy (ACC) and average backward transfer (BWT) across $5$ independent runs with five seeds under the same experimental setup [17]. $\dagger$    denotes results reported from [17].